# TEAL: New Selection Strategy for Small Buffers in Experience Replay Class Incremenal Learning

### v1

@@ 4.3 Ablation study
We explore an alternative variant of TEAL, which selects the set of exemplars in a single pass. We call this variant TEAL_OneTime. While still selecting exemplars for each class separately, this variant first partitions the exemplars of each class into $n=\frac{|{\mathcal{M}}|}{N^{t}}$ clusters, and then selects the most typical exemplar from each cluster in descending order of cluster size, preserving the order of selections.
The experiments on the two variants are conducted using Split CIFAR-100 with the smaller version of ResNet-18 mentioned in Section 4.1 and buffer sizes of 300, 500 and 2000. We investigate the performance enhancement of each variant when integrated with XDER, ER-ACE, and ER. The results are shown in Table 2. Whenever there is a significant difference in performance, the iterative variant outperforms the other one in 4 out of 5 cases. However, the differences are not substantial.
@@ 5 Summary and discussion
We propose a new mechanism to select exemplars for the memory buffer in replay-based CIL methods. This method is based on the principles of diversity and representativeness. When the memory buffer is relatively small, our method TEAL is shown to outperform both the native mechanism of each ER-IL method (usually random selection) and an alternative selection mechanism called Herding. Even when the buffer is large, our method is beneficial in almost all studied cases.
Our future work will investigate ways to determine at which point, if any, the added value of our method diminishes or possibly becomes harmful. We will also integrate TEAL with other replay-based CIL methods that utilize different selection strategies, and analyze the impact on their performance.
Acknowledgement
This work was supported by grants from the Israeli Council of Higher Education and the Gatsby Charitable Foundations.
Appendix
@@ Appendix A Implementation details
The source code used in this study would be uploaded to our GitHub repository and will be made public upon acceptance.
@@
A.1 Stand-alone setting
@@
  1. 1.
    The Order we display on Fig. 5 is: [44, 19, 93, 90, 71, 69, 37, 95, 53, 91, 81, 42, 80, 85, 74, 56, 76, 63, 82, 40, 26, 92, 57, 10, 16, 66, 89, 41, 97, 8, 31, 24, 35, 30, 65, 7, 98, 23, 20, 29, 78, 61, 94, 15, 4, 52, 59, 5, 54, 46, 3, 28, 2, 70, 6, 60, 49, 68, 55, 72, 79, 77, 45, 1, 32, 34, 11, 0, 22, 12, 87, 50, 25, 47, 36, 96, 9, 83, 62, 84, 18, 17, 75, 67, 13, 48, 39, 21, 64, 88, 38, 27, 14, 73, 33, 58, 86, 43, 99, 51]
  2. 2.
    The order in 7(a) is: [45, 15, 90, 32, 35, 63, 17, 72, 79, 96, 48, 36, 16, 11, 23, 80, 22, 58, 3, 62, 50, 33, 66, 99, 43, 76, 7, 57, 81, 82, 6, 10, 24, 52, 95, 73, 91, 21, 38, 31, 85, 59, 13, 69, 75, 70, 64, 8, 77, 34, 46, 39, 92, 0, 44, 98, 49, 9, 4, 61, 12, 83, 28, 78, 40, 88, 54, 5, 26, 41, 89, 20, 84, 2, 1, 55, 19, 74, 25, 37, 42, 14, 30, 18, 67, 71, 68, 27, 60, 51, 29, 56, 93, 47, 97, 94, 86, 87, 65, 53]
  3. 3.
    The order in 7(b) is: [48, 97, 1, 81, 90, 49, 10, 8, 7, 20, 70, 73, 75, 14, 91, 38, 47, 21, 74, 52, 80, 98, 59, 12, 71, 85, 6, 34, 55, 82, 95, 63, 78, 15, 94, 60, 99, 76, 25, 40, 88, 0, 62, 96, 87, 51, 16, 18, 9, 19, 29, 45, 86, 53, 56, 31, 28, 61, 30, 33, 4, 67, 64, 58, 50, 54, 3, 13, 37, 27, 66, 77, 84, 69, 2, 41, 22, 92, 42, 44, 11, 36, 46, 79, 65, 72, 23, 17, 39, 5, 89, 35, 24, 83, 43, 57, 93, 32, 68, 26]
@@
Figure 7: Enhanced accuracy of TEAL compared to random-sampling and Herding selection strategies, additional two different classes order.
A.2 Integrated setting
Here we train a ResNet-18 model for 100 epochs using the same optimizer as in A.1, with the same batch size and data augmentations, with some exceptions. ER-ACE starts with a learning rate of 0.01, and all experiments on Split CUB-200 are conducted for 30 epochs with a batch size of 16 due to the dataset size and the resolution of its images.
A.3 Baselines setting
The setting for the experiments of the baseline methods is the same as in A.2 with some exceptions. For BiC, the number of training epochs is 250, and the learning rate scheduler decays the learning rate by a factor of 0.1 on epochs 100, 150 and 200. For GEM, the batch size is 32 and the learning rate starts from 0.03.For GDumb, the batch size is 32. For iCaRL the learning rate starts from 2, the weight decay is 0.00001 and the learning rate scheduler decays the learning rate by a factor of 0.2 on epochs 49 and 63.
Appendix B Compute resources

### v2

@@ 4.3 Ablation study
Iterative process of TEAL
We explore an alternative variant of TEAL, which selects the set of exemplars in a single pass instead of by iterations. We call this variant TEAL_OneTime. While still selecting exemplars for each class separately, this variant first partitions the exemplars of each class into $n=\frac{|{\mathcal{M}}|}{N^{t}}$ clusters, and then selects the most typical exemplar from each cluster in descending order of cluster size, preserving the order of selections.
Table 3: The improvement of 2 variants of TEAL when integrated with different CIL methods: XDER, ER-ACE, and ER. In almost all cases, the original TEAL matches or outperforms the OneTime variant.
$|{\mathcal{M}}|$ Selection Strategy XDER ER-ACE ER
300 TEAL_OneTime 44.13±0.26 35.81±0.42 17.64±0.12
TEAL 45.05±0.24 36.4±0.12 18.1±0.1
500 TEAL_OneTime 49.87±0.22 40.16±0.3 22.01±0.18
TEAL 50.29±0.2 40.26±0.23 22.07±0.13
2000 TEAL_OneTime 59.05±0.14 51.52±0.24 40.65±0.32
TEAL 58.79±0.17 51.68±0.35 39.96±0.23
Table 4: Final average accuracy for method with and without TEAL using the ArchCraft architecture. Top 3 rows correspond to Split CIFAR-100, bottom 3 to Split tinyImageNet.
ER-ACE ER
$|{\mathcal{M}}|$ Vanilla +TEAL Imp. Vanilla +TEAL Imp.
300 36.75±0.28 42.47±0.08 5.72 14.59±0.08 18.01±0.35 3.42
500 41.98±0.24 47.53±0.35 5.56 18.54±0.4 22.3±0.4 3.76
2000 57.33±0.18 60.18±0.17 2.85 41.03±0.37 43.44±0.28 2.41
400 12.24±0.11 13.2±0.38 0.96 7.67±0.13 7.84±0.19 0.17
1000 15.64±0.37 16.7±0.2 1.07 8.12±0.05 8.36±0.18 0.24
2000 19.08±0.36 20.43±0.37 1.35 9.65±0.11 10.66±0.03 1.01
The experiments on the two variants are conducted using Split CIFAR-100 with the smaller version of ResNet-18 mentioned in Section 3 and buffer sizes of 300, 500 and 2000. We investigate the performance enhancement of each variant when integrated with XDER, ER-ACE, and ER. The results are shown in Table 3. Clearly, whenever there is a significant difference in performance, the iterative variant outperforms the other one in 4 out of 5 cases.
Different architecture
Some experiments were replicated using ArchCraft instead of the ResNet-18 architecture, obtaining similar results (see Table 4).
@@ 5 Summary and discussion
We proposed a new mechanism to select exemplars for the memory buffer in replay-based CIL methods. This method is based on the principles of diversity and representativeness. When the memory buffer is relatively small, our method TEAL is shown to outperform both the native mechanism of each ER-IL method (usually random selection) and alternative selection mechanisms, including Herding, Uncertainty and Centered. Even when the buffer is large, our method is beneficial in almost all cases.
In future work we will investigate ways to determine at which point, if any, the added value of our method diminishes or possibly becomes harmful. We will also integrate TEAL with other replay-based CIL methods that utilize different selection strategies.
Supplementary Material
@@ Appendix A Implementation details
The source code for this study can be found in the zip file included in the supplementary material. This zip file contains a README.md file that provides instructions on how to run the experiments. The code would be uploaded to our GitHub repository and will be made public upon acceptance.
A.1 TEAL implementation
Clustering algorithm
We used scikit-learn KMeans implementation.
Iterations pace
@@ A.1 TEAL implementation
A.2 Stand-alone setting
@@
1. The order we display on Fig. 7 is: [44, 19, 93, 90, 71, 69, 37, 95, 53, 91, 81, 42, 80, 85, 74, 56, 76, 63, 82, 40, 26, 92, 57, 10, 16, 66, 89, 41, 97, 8, 31, 24, 35, 30, 65, 7, 98, 23, 20, 29, 78, 61, 94, 15, 4, 52, 59, 5, 54, 46, 3, 28, 2, 70, 6, 60, 49, 68, 55, 72, 79, 77, 45, 1, 32, 34, 11, 0, 22, 12, 87, 50, 25, 47, 36, 96, 9, 83, 62, 84, 18, 17, 75, 67, 13, 48, 39, 21, 64, 88, 38, 27, 14, 73, 33, 58, 86, 43, 99, 51]
2. The order in Fig. 8(a) is: [45, 15, 90, 32, 35, 63, 17, 72, 79, 96, 48, 36, 16, 11, 23, 80, 22, 58, 3, 62, 50, 33, 66, 99, 43, 76, 7, 57, 81, 82, 6, 10, 24, 52, 95, 73, 91, 21, 38, 31, 85, 59, 13, 69, 75, 70, 64, 8, 77, 34, 46, 39, 92, 0, 44, 98, 49, 9, 4, 61, 12, 83, 28, 78, 40, 88, 54, 5, 26, 41, 89, 20, 84, 2, 1, 55, 19, 74, 25, 37, 42, 14, 30, 18, 67, 71, 68, 27, 60, 51, 29, 56, 93, 47, 97, 94, 86, 87, 65, 53]
3. The order in Fig. 8(b) is: [48, 97, 1, 81, 90, 49, 10, 8, 7, 20, 70, 73, 75, 14, 91, 38, 47, 21, 74, 52, 80, 98, 59, 12, 71, 85, 6, 34, 55, 82, 95, 63, 78, 15, 94, 60, 99, 76, 25, 40, 88, 0, 62, 96, 87, 51, 16, 18, 9, 19, 29, 45, 86, 53, 56, 31, 28, 61, 30, 33, 4, 67, 64, 58, 50, 54, 3, 13, 37, 27, 66, 77, 84, 69, 2, 41, 22, 92, 42, 44, 11, 36, 46, 79, 65, 72, 23, 17, 39, 5, 89, 35, 24, 83, 43, 57, 93, 32, 68, 26]
@@
Figure 8: Split CIFAR-100: TEAL Enhanced accuracy of TEAL compared to 4 baselines: random-sampling, Herding, Centered, and Uncertainty, in two different classes order.
A.3 Integrated setting
Here we train a ResNet-18 model for 100 epochs using the same optimizer as in A.2, with the same batch size and data augmentations, with some exceptions. ER-ACE starts with a learning rate of 0.01 and train on batch size of 10 examples as in the original paper, and all experiments on Split CUB-200 are conducted for 30 epochs with a batch size of 16 due to the dataset size and the resolution of its images. In experiments using the ArchCraft model, we used the ResAC-A model as implemented in the official paper’s code.
A.4 Baselines setting
The setting for the experiments of the baseline methods is the same as in A.3 with some exceptions. For BiC, the number of training epochs is 250, and the learning rate scheduler decays the learning rate by a factor of 0.1 on epochs 100, 150 and 200. For GEM, the batch size is 32 and the learning rate starts from 0.03.For GDumb, the batch size is 32. For iCaRL the learning rate starts from 2, the weight decay is 0.00001 and the learning rate scheduler decays the learning rate by a factor of 0.2 on epochs 49 and 63.
A.5 Compute resources
@@
Appendix B Additional Results
(a) Split CIFAR-100
(b) Split tinyImageNet
Figure 9: Performance improvement of TEAL when integrated with ER and with ER-ACE over various buffer sizes.
(a) Split CIFAR-100, $|{\mathcal{M}}|=500$
(b) Split CIFAR-100, $|{\mathcal{M}}|=2000$
(c) Split tinyImageNet, $|{\mathcal{M}}|=1000$
(d) Split tinyImageNet, $|{\mathcal{M}}|=2000$
(e) Split CUB-200, $|{\mathcal{M}}|=400$
Figure 10: The performance of ER-IL methods with and without TEAL. First row displays the average accuracy after training incrementally on a different number of classes. Each color corresponds to a different ER-IL method, where the continuous line represents the vanilla method, while the dashed line represents the method with TEAL as its selection strategy. The error bars correspond to standard error based on 4-10 repetitions. Second row depicts the difference in accuracy between TEAL and another method (XDER, ER-ACE, and ER) across all tasks.
(a) $|{\mathcal{M}}|=300$
(b) $|{\mathcal{M}}|=500$
(c) $|{\mathcal{M}}|=2000$
Figure 11: Split CIFAR-100: The performance of ER-IL methods with and without TEAL with splitting CIFAR-100 into 20 tasks instead of 10. Each color corresponds to a different ER-IL method, where the continuous line represents the vanilla method, while the dashed line represents the method with TEAL as its selection strategy. The error bars correspond to standard error based on 6 repetitions.
(a) $|{\mathcal{M}}|=500$
(b) $|{\mathcal{M}}|=2000$